### v1

# A Fully Automated Pipeline for Conversational Discourse Annotation: Tree Scheme Generation and Labeling with Large Language Models

Abstract
Recent advances in Large Language Models (LLMs) have shown promise in automating discourse annotation for conversations. While manually designing tree annotation schemes significantly improves annotation quality for humans and models, their creation remains time-consuming and requires expert knowledge. We propose a fully automated pipeline that uses LLMs to construct such schemes and perform annotation. We evaluate our approach on speech functions (SFs) and the Switchboard-DAMSL (SWBD-DAMSL) taxonomies. Our experiments compare various design choices, and we show that frequency-guided decision trees, paired with an advanced LLM for annotation, can outperform previously manually designed trees and even match or surpass human annotators while significantly reducing the time required for annotation. We release all code and resultant schemes and annotations to facilitate future research on discourse annotation.
1 Introduction
Discourse analysis is essential in NLP tasks like dialog management, generation, summarization, and emotion recognition (Liang et al., 2020; Chen et al., 2021; Shou et al., 2022). Traditionally, discourse annotation depends on manual expert labeling, which is costly and time-consuming. LLM-based annotation presents a promising alternative, enhancing speed, consistency, and cost-effectiveness (Gilardi et al., 2023; Hao et al., 2024). However, challenges such as biases and domain limitations necessitate careful prompt design and continuous evaluation.
Figure 1: Example of speech function structure Ostyakova et al. (2023).
In Ostyakova et al. (2023), the authors explored using ChatGPT to automate discourse annotation for labeling chit-chat dialogs using the speech functions (SFs) taxonomy (Eggins and Slade, 2004). SFs categorize communicative acts in dialog, capturing speaker intentions and interactions in a hierarchical structure to analyze conversational flow (see Figure 1 for an example and Appendix A for the full label set). Ostyakova et al. (2023) conducted three sets of experiments: (1) Direct Annotation: The LLM assigned labels from a predefined list of SFs; (2) Step-by-Step Scheme with Intermediate Labels: LLM selected labels progressively from broad to specific categories; (3) Complex Tree-Like Scheme with Yes/No Questions: using a complex tree-like annotation pipeline originally designed for crowdsourced annotation. Since breaking a multi-label selection task into smaller sub-tasks using a tree structure has improved human performance in complex discourse annotation (Scholman et al., 2016), the authors hypothesized that the same approach could enhance LLM-based annotation. Prior research also suggests that guiding models with tree-structured prompts significantly improves performance (Yao et al., 2024).
Ostyakova et al. (2023) found that the Tree-Like Scheme approach enhances LLM accuracy, achieving near-human performance, and suggest that LLMs could serve as a “silver standard” for annotation. However, newer and more powerful LLMs have been released since this work was published. These developments not only enhance the potential of LLMs to be used for annotation but also open the door to automating the creation of tree-like schemes, [^1] We will refer to them as tree schemes. enabling a fully automated pipeline. This is especially valuable for large taxonomies, such as Intelligent PAL (Morrison et al., 2014) in the educational domain and TEACh-DA (Gella et al., 2022) for task-oriented systems, where manually creating tree schemes is complex and time-intensive.
This work automates tree schemes creation, making them usable for annotation by crowdsourced workers and LLMs. A tree scheme is a decision tree that classifies dialog utterances through a series of questions, which can be binary or non-binary, using yes/no or open-ended formats. An example of a tree generated using the pipeline proposed in this work can be found in Appendix B. We provide a repository with the code and generated trees to support further research and applications. [^2] https://github.com/Kpetyxova/autoTree
2 Related Work
Discourse Analysis
Researchers analyze discourse structures to improve dialog understanding and management, focusing on pragmatics and speaker intent. Two key frameworks are the Dialog Act (DA) Theory Jurafsky et al. (1998), which assigns pragmatic labels to utterances, and Segmented Discourse Representation Theory (SDRT) Lascarides and Asher (2007), which models relationships between discourse units. [^3] These frameworks are presented as general background and are not used in the experiments.
The SWBD-DAMSL scheme Jurafsky et al. (1997a), initially created for casual conversations and widely applied to task-oriented systems, classifies dialog acts into 42 classes. Taxonomy of speech functions Eggins and Slade (2004) offers a hierarchical annotation approach, integrating DA principles and relational analysis.
Dialog acts are beneficial in task-oriented dialog agents. For example, Gella et al. (2022) introduce a scheme for embodied agents, improving natural language interactions and task success, and Leech and Weisser (2003) develop Speech Act Annotated Corpus (SPAAC) scheme, balancing specificity and generalizability in task-oriented dialogs.
LLMs for Discourse and Annotation
Ostyakova et al. (2023) introduce a semi-automated approach for annotating open-domain dialogs using a taxonomy of speech functions and ChatGPT. Their study evaluates three methods: Direct Annotation (selecting from a complete label set), Step-by-Step (progressively narrowing choices), and Tree-Like Schemes (using hierarchical yes/no questions). The Tree-Like Scheme has performed best, particularly for rare classes, achieving high consistency when running the same annotation pipeline with ChatGPT three times (Fleiss’ kappa: 0.83). However, expert input has remained essential for designing the annotation pipelines.
Yadav et al. (2024) and Tseng et al. (2024) explore GPT-4-based semantic annotation, emphasizing prompt design and model limitations. Chen et al. (2024) investigate LLMs for event extraction, addressing data scarcity in fine-tuned models. Finally, Wu et al. (2024) introduce a rationale-driven collaborative framework that refines annotations through iterative reasoning, outperforming standard prompting.
3 Speech Functions Corpus
The experiments described in this work are based on the Speech Functions Corpus, a dataset of dialogs annotated with SFs. This corpus was developed by Ostyakova et al. (2023), where three experts, each with at least a B.A. in Linguistics, annotated the DailyDialog dataset (Li et al., 2017) – a multi-turn casual dialog dataset – using the speech functions taxonomy (Eggins and Slade, 2004). The authors of the corpus reduced the original 45 classes proposed by Eggins and Slade (2004) to a more manageable set of 32 classes.
The tag set covers five functional dimensions: turn management, topic organization, feedback, communicative acts, and pragmatic purposes. While all dimensions are embedded within speech functions, they are distributed unevenly across tags, with individual speech functions incorporating between two and five dimensions. Figure 1 illustrates an example of a speech function that includes all five dimensions: React.Rejoinder.Support.Track.Clarify. In this example: (1) React represents turn management, indicating a speaker change or a reaction to a previous utterance; (2) Rejoinder corresponds to topic organization, signifying active topic development that influences the dialog flow; (3) Support denotes feedback, showing that the speaker is supporting an interlocutor; (4) Track falls under communicative acts, identifying questions; (5) Clarify serves a pragmatic purpose, indicating a question aimed at obtaining additional information on the current conversation topic. Some SFs, however, cover fewer dimensions. For instance, Open.Attend represents only two: turn management (marking the conversation’s beginning) and communicative acts (a greeting).
The Speech Functions Corpus includes 64 dialogs, containing 1,030 utterances. Appendix C shows an example of an annotated dialog.
To evaluate the effect of the taxonomy size on the method proposed in this work, the existing taxonomy was converted into the following subsets:
  • Full Taxonomy: This set includes all 32 speech functions labels from Ostyakova et al. (2023). A complete list of labels, along with their descriptions, examples, and frequency information, can be found in Appendix A.
  • The Top Level of the Taxonomy: This subset consists of three classes corresponding to the turn management level of the speech functions taxonomy. Definitions for these labels were written manually and are detailed in Appendix D. This subset is particularly important because if the model fails to distinguish between these high-level categories, the entire hierarchical structure may be unreliable.
  • The Top Two Levels of the Taxonomy: This subset includes six classes, representing a combination of the speech functions taxonomy’s turn management and topic organization dimensions. Manual definitions for these classes are provided in Appendix E. The motivation for analyzing this subset is similar to the previous one but with an additional level of complexity, making the classification task more challenging.
  • Top-20 Frequent Classes of the Taxonomy: This subset comprises the 20 most frequently occurring classes in the Speech Functions Corpus. It is designed to evaluate how well the model handles frequent classes in the absence of rarer ones.
4 Framework for Tree Construction with LLMs
Traditionally, discourse annotation has been performed manually by experts or trained annotators, relying on predefined end-class descriptions. However, Scholman et al. (2016) investigates whether non-trained, non-expert annotators can reliably annotate coherence relations using a step-wise approach, which functions similarly to a decision tree. Their findings indicate that a structured step-wise method can indeed make discourse annotation more accessible to non-experts, facilitating large-scale annotation without requiring extensive training. This is achieved by using cognitively plausible primitives rather than relying on complex end labels. A similar observation is made in Ostyakova et al. (2023), further supporting the viability of this approach.
To construct an effective decision tree for annotation, it is essential to design questions that do not require expert-level knowledge of discourse but can instead be answered based on the utterance itself. Ideally, related classes should be positioned closely within the tree, forming a hierarchical structure that reflects conceptual similarities between coherence relations. This hierarchical organization not only simplifies decision-making for annotators but also enhances consistency and reliability in annotation.
Tree Construction
The pipeline for the tree construction process is illustrated in Figure 2. The core concept of the proposed algorithm is to use an advanced LLM to identify distinguishing features that can divide classes into two or more groups. The LLM is provided with a table containing class names, definitions, and usage examples as input.
Figure 2: Pipeline for tree construction. LLM formulates a classification question to split classes into groups, mapping possible answers to respective class groups. This process repeats recursively for created groups until all groups contain only one class. Finally, the grouped data is merged into a single tree structure in JSON format for annotation.
To enhance reasoning capabilities, we aim for the model to engage in an inner monologue (Zhou et al., 2024) before determining how to split the data. To achieve this, the LLM is prompted to first generate a set of questions about an utterance that can aid its understanding. Specifically, the model is instructed to formulate and answer three such reasoning question-answer pairs at the beginning of its response. Recent studies indicate that employing such self-questioning techniques helps models produce more flexible, meaningful outputs, indicating a deeper level of comprehension and reasoning (Sun et al., 2024). An example of the LLM’s reasoning output can be found in Appendix F.
After this reasoning step, the LLM generates a classification question to determine an utterance’s group, providing possible answers mapped to class groups. For example, the model might generate a classification question: “Is this the beginning of a conversation?” The possible answers (groups) could be: (1) “Yes, this is the beginning of a conversation.” and (2) “No, the utterance continues the conversation.” This question helps categorize utterances, splitting the taxonomy into conversation openings and other categories. The prompt template used for this process is detailed in Appendix G.
With taxonomy classes grouped, the process iterates through each group until no group contains more than one class. Once resolved, the resulting JSONs are merged into a single tree JSON for annotation. For this step, GPT-4 (gpt-4-0613) is used due to its ability to handle long contexts and generate valid JSON outputs OpenAI (2023). The temperature is set to 0.4, which is empirically chosen to balance creativity and reliability, allowing some variation while avoiding excessive randomness. In contrast, a temperature of 0 minimizes variability but may produce overly rigid outputs.
Annotation
Using the constructed questions tree, dialogs are annotated by traversing the tree for each utterance until reaching a leaf node representing a class label. To ensure a fair comparison with the manually created tree from Ostyakova et al. (2023), the same annotation settings are applied. Specifically, the context length is set to 1, as this yielded the best results in ablation studies. The same model – ChatGPT (GPT-3.5-turbo) – is used, along with the optimal temperature of 0.9, as identified in the ablation studies. The prompt format remains consistent with Ostyakova et al. (2023) and is shown in Appendix H. Additionally, open-source models’ performance is analyzed in Section 6.
Evaluation
The framework’s performance is assessed by comparing the predicted labels from the annotation step to the gold references provided in Ostyakova et al. (2023), using the same 12 dialogs (189 utterances) as the authors, referred to here as the development set. To demonstrate generalizability across dialogs, we also evaluate it on another 12 dialogs (165 utterances) from the Speech Functions corpus, referred to as the test set. To further test generalizability across datasets, we evaluate it on a subset of the Switchboard Dialog Act Corpus — a dataset of telephone conversations annotated using the SWBD-DAMSL annotation scheme, as described in Section 7. The evaluation uses the same metrics as in Ostyakova et al. (2023): Weighted Precision ($P_{w}$), Weighted Recall ($R_{w}$), and Macro F1 ($F1$). Additionally, we report Weighted F1 ($F1_{w}$).
5 Experiments & Results
5.1 Yes/No Questions vs. Open-Ended Questions
The manually created tree in Ostyakova et al. (2023) is binary and composed exclusively of yes/no questions. In this subsection, we implement a binary yes/no question tree using the proposed framework. Additionally, we construct a binary tree with open-ended questions. The hypothesis behind this experiment is that allowing the model to generate any questions, rather than limiting it to yes/no questions, can result in a more flexible and nuanced tree for discourse annotation. The prompts used for constructing both the yes/no questions and the open-ended questions tree were developed through iterative refinement. This refinement process involved adjusting and testing prompts until they produced stable and consistent results, ensuring a valid JSON output with all necessary keys present. The final versions of these prompts are provided in Appendix I and Appendix J, respectively.
Evaluation results for the Top Level, the Top Two Levels, and Top-20 class subsets on the development set are presented in Tables 12 and 3 while the results for the complete SF taxonomy on the development and test sets are shown in Tables 4 and 5, respectively. [^4] The best results for GPT-3.5 are highlighted in bold, while the second-best results are marked with an underline. The overall best results across all models are highlighted in blue. The findings indicate that open-ended trees outperform yes/no trees across all metrics and data subsets. However, for the complete SF taxonomy, both yes/no and open-ended trees perform significantly worse than the manually created tree.
Approach $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
Yes/no 0.55 0.34 0.37 0.33
Open-ended 0.68 0.63 0.61 0.63
Non-binary 0.74 0.72 0.70 0.73
Table 1: Evaluation of annotations on the development set using trees constructed through different methods for the Top Level of the SF taxonomy, with GPT-3.5 used for annotation.
Approach $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
Yes/no 0.49 0.22 0.26 0.20
Open-ended 0.70 0.65 0.65 0.43
Non-binary 0.60 0.48 0.45 0.48
W/ split selection 0.67 0.62 0.62 0.60
Freq.-guided split selection 0.63 0.45 0.41 0.55
[…]
Freq.-guided split selection (GPT-4o for annot.) 0.80 0.78 0.78 0.80
Table 2: Evaluation of annotations on the development set using trees constructed through different methods for the Top Two Levels of the SF taxonomy (with GPT-3.5 used for annotation unless explicitly stated otherwise).
Approach $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
Yes/no 0.36 0.18 0.16 0.14
Open-ended 0.42 0.40 0.37 0.19
Non-binary 0.51 0.26 0.22 0.20
W/ split selection 0.39 0.36 0.34 0.27
Freq.-guided split selection 0.62 0.66 0.62 0.35
[…]
Freq.-guided split selection (GPT-4o for annot.) 0.70 0.69 0.67 0.41
Table 3: Evaluation of annotations on the development set using trees constructed through different methods for the Top-20 classes of the SF taxonomy (with GPT-3.5 used for annotation unless explicitly stated otherwise).
Approach $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
Manually created tree from Ostyakova et al. (2023) (crowdsourced annotation of the full dataset) 0.71 0.60 - 0.46
Manually created tree from Ostyakova et al. (2023) (ChatGPT for annotation) 0.67 0.62 - 0.43
Yes/no 0.37 0.25 0.24 0.13
Open-ended 0.38 0.23 0.21 0.23
Non-binary 0.39 0.34 0.31 0.16
W/ split selection 0.36 0.38 0.35 0.16
Freq.-guided split selection 0.31 0.43 0.34 0.19
W/ split selection (GPT-4o for annotation) 0.57 0.53 0.51 0.32
Freq.-guided split selection (GPT-4o for annotation) 0.83 0.75 0.74 0.60
W/ split selection (Llama-3.1-8B-Instruct for annotation) 0.56 0.40 0.41 0.24
Freq.-guided split selection (Llama-3.1-8B-Instruct for annotation) 0.45 0.48 0.41 0.30
W/ split selection (Mistral-7B-Instruct-v0.3 for annotation) 0.48 0.50 0.54 0.31
Freq.-guided split selection (Mistral-7B-Instruct-v0.3 for annotation) 0.33 0.44 0.32 0.18
Table 4: Evaluation of annotations on the development set (except for the first line) using trees constructed through different methods (with GPT-3.5 used for annotation unless explicitly stated otherwise).
Approach $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
Yes/no 0.20 0.21 0.17 0.12
Open-ended 0.35 0.22 0.21 0.16
Non-binary 0.31 0.31 0.27 0.16
W/ split selection 0.48 0.18 0.16 0.17
Freq.-guided split selection 0.43 0.42 0.37 0.23
[…]
Table 5: Evaluation of annotations on the test set using trees constructed through different methods (with GPT-3.5 used for annotation unless explicitly stated otherwise).
5.2 Binary vs. Non-Binary Trees
Allowing the model to split nodes into multiple groups, rather than restricting it to two groups, has the potential to create a more granular and detailed annotation tree. However, this added granularity may introduce greater complexity and reduce consistency in the annotation process. [^5] This experiment is distinct from the previous one, where we only allowed binary open-ended questions and compared them to yes/no questions. As a result, both the questions and answers in that setting were more high-level, distinguishing data based on a single characteristic. In contrast, non-binary open-ended questions allow for greater granularity and specificity, enabling more nuanced differentiation within the data. Experiments described in this section examine the impact of restricting the model to binary splits compared to allowing multiple-group splits on the accuracy of the resulting annotation tree. Both the binary and non-binary trees discussed in this section use open-ended questions. A prompt that allows open-ended questions and splitting data into more than two groups is detailed in Appendix K.
The findings indicate that allowing the model to split data into multiple groups generally outperforms restricting it to binary splits (Tables 1, 2, 3, 4 and 5). This performance difference is more pronounced when the number of classes is smaller. However, as the number of classes increases, the gap narrows, with weighted metrics occasionally favoring the binary approach, specifically showing (1) higher $P_{w}$ and $R_{w}$ for the Top Two Levels (6 classes); (2) higher $R_{w}$ for the Top-20 classes; (3) higher $P_{w}$ for the complete taxonomy (33 classes) on the test set. Nevertheless, higher macro metrics for non-binary trees suggest improved performance for smaller and less frequent classes. Based on these results, further experiments will allow the model to split data into more than two groups.
5.3 Optimal Split Selection and Backtracking
Inspired by Yao et al. (2024), we evaluate the impact of allowing the model to backtrack while constructing the annotation tree. In each iteration, the model generates three potential splits and assigns a score to each (the prompt used for scoring is provided in Appendix L). These splits are then evaluated using a pre-trained natural language inference (NLI) model, [^6] https://huggingface.co/cross-encoder/nli-deberta-v3-base which classifies them as either contradictory, neutral, or non-contradictory. Among the non-contradictory options, the split with the highest score is selected. If the best-scoring split does not produce a viable partition, the model backtracks and evaluates the next-best option. The motivation behind this experiment is that multiple valid ways to create splits exist, and selecting the seemingly best option at each step may not always result in the most effective tree overall.
For the Two-Level subset, performance improves compared to the approach that does not use split selection and backtracking (Table 2). In the Top-20 subset, only the $F1$ metric shows an increase, indicating better performance for less frequent classes (Table 3). For the complete SF taxonomy, performance on the development set remains comparable to the approach without split selection and backtracking (Table 4), while on the test set, the $F1$ score is higher (Table 5). [^7] Results for the Top Level subset are unavailable, as these trees have only one level.
The lack of improvement in the complete SF taxonomy stems from the absence of an optimal split at the initial step, causing error propagation throughout the taxonomy. Specific issues include: (1) The model misassigned the Open.Attend category (which represents greetings at the beginning of a conversation) to the branch “The utterance involves a request for information,” and (2) It grouped all React classes under the branch “The dialog utterance involves a response to a request for information,” which does not accurately represent them. This misclassification often led the annotation model to misroute utterances to the Sustain branch instead, resulting in unreliable annotations. These issues are not due to the updated approach but rather to the fundamental challenge of generating meaningful splits when dealing with many classes. These errors propagate throughout the taxonomy if the model fails to establish a strong initial split.
5.4 Frequency-Guided Optimal Split Selection and Backtracking
Class frequency information can be used to guide the model in making splits and optimize the annotation process. In conversations, certain SFs occur more frequently than others. For instance, Sustain.Continue.Prolong.Extend appears 21.8% of the time, representing instances where a speaker adds information to their preceding statement. Similarly, React.Rejoinder.Support.Track.Clarify occurs 12% of the time, typically indicating a question aimed at obtaining additional information. Meanwhile, some SFs are relatively rare. The experiments described in this section aim to construct a decision tree that reflects the distribution of classes, making frequent classes easier and faster to reach compared to rare ones.
To achieve a frequency-guided tree, the prompt used to generate splits is modified as follows: at each step, the model is instructed to create a group containing only the most frequent class if one class is significantly more frequent than the others. The full prompt is provided in Appendix M.
The results indicate that the metrics for the Two-Level subset have decreased compared to the approach without frequency guidance (Table 2), while they have significantly increased for the Top-20 subset (Table 3). For the complete SF taxonomy, the metrics remained at the same level on the development set (Table 4) but improved on the test set (Table 5).
Manual analysis revealed that during the annotation step, the model frequently selected incorrect paths, often defaulting to upper-level classes. This behavior was especially prevalent in the most frequent class, Sustain.Continue.Prolong.Extend. Despite this, the tree structure appears logical, and the root question is straightforward: Does the dialog utterance provide supplementary or contradictory information to the previous statement by the same speaker? The response options for this question are: (1) Dialog utterances that provide supplementary or contradictory information to the previous statement by the same speaker; (2) Dialog utterances that do not provide supplementary or contradictory information to the previous statement by the same speaker. While the question is specific, emphasizing conditions about the same speaker and the addition of information, the model often ignored these requirements. In numerous cases, the first response option was incorrectly assigned, even at the start of a conversation.
The distinguishing characteristic of this tree is the heightened granularity and specificity of the questions and labels, with each step designed to determine whether the utterance fits a particular class using a single, targeted question. To assess whether the frequency-guided tree presents too significant a challenge for the GPT-3.5 model and whether it might perform better with a more advanced model, GPT-4o Hurst et al. (2024) was used during the annotation step (see Section 5.4.1). This approach allowed for a direct comparison between GPT-4o and GPT-3.5. Section 5.4.2 also examines whether the observed differences in metrics are the same for trees created without frequency guidance.
5.4.1 GPT-4o for Annotation
Evaluation results comparing annotations by GPT-4o and GPT-3.5 on trees constructed using frequency-guided optimal split selection and backtracking for the Two-Level and Top-20 class subsets are presented in Tables 2 and 3. Results for the complete SF taxonomy on the development and test sets are shown in Tables 4 and 5, respectively. We note that not only are the differences in metrics highly significant, but also the $\mathbf{P_{w}}$, $\mathbf{R_{w}}$, and $\mathbf{F1}$ scores for annotations on both the development and test sets for the complete SF taxonomy surpass those obtained for the entire dataset annotated by crowdsourcers in Ostyakova et al. (2023). This finding underscores that the proposed Frequency-Guided Optimal Split Selection approach, combined with an advanced LLM for annotation, may both outperform manually constructed trees and improve traditional human-driven annotation processes.
5.4.2 Do Annotation Gaps Persist in Non-Frequency-Guided Approaches?
This section examines whether the significant differences observed between annotations generated by GPT-4o and GPT-3.5 also occur in the non-frequency-guided optimal split selection and backtracking approach.
Tables 234 and 5 presents performance metrics for annotations produced using GPT-3.5 and GPT-4o with the optimal split selection algorithm, both with and without frequency guidance. For the frequency-guided approach, the performance gap between GPT-3.5 and GPT-4o becomes more pronounced as the number of classes increases. In contrast, there is no significant difference in performance for the non-frequency-guided approach when switching from GPT-3.5 to GPT-4o. This underscores the clear advantage of the frequency-guided approach when paired with a more advanced model.
5.5 Cost Analysis
This section provides a cost analysis for creating trees for SFs and annotating data using these trees.
Generating a non-binary tree for the full SF taxonomy with GPT-4 costs approximately \$0.40 and takes 2 minutes (max depth: 3). Annotation with GPT-4o costs \$0.24 per dialog, taking 50 seconds.
Using frequency-guided optimal split selection and backtracking, tree creation costs \$5.48 (with \$4.05 for split-candidates and \$1.43 for scoring) and takes 32 minutes. Without optimal split selection and backtracking, a frequency-guided tree costs \$1.83. In this case, the maximum tree depth is around 7. Annotation costs approximately \$0.36 per dialog and takes about 35 seconds.
For comparison, GPT-3.5 annotation with a manually created tree costs \$0.03–\$0.07 per dialog Ostyakova et al. (2023). Crowdsourced annotation costs \$0.12–\$0.22 per dialog, averaging 29 minutes per annotation. While the authors do not specify the time required for tree creation, assuming it exceeds half an hour is reasonable.
Based on these estimates, annotating the entire dataset (64 dialogs) using only human resources would cost approximately \$10.88 and take 31 hours plus additional time for tree creation. In contrast, our best approach, frequency-guided optimal split selection with GPT-4o, would cost around \$20.84 but reduce the total time to approximately 1 hour and 25 minutes, offering significant efficiency and quality benefits despite the higher cost.
6 Open-Source Models for Annotation
Table 4 presents the results of using two open-source models, Mistral-7B-Instruct-v0.3888https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct, [^9] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct compared to the closed-source models GPT-3.5 and GPT-4o for the annotation steps with two approaches, frequency-guided and non-frequency-guided optimal split selection with backtracking, which were selected based on their strong performance with other models.
The results indicate that despite having fewer parameters than the closed-source models, both open-source models significantly outperform GPT-3.5. Specifically, when using the non-frequency-guided approach, both open-source models achieve performance close to GPT-4o while markedly surpassing GPT-3.5. However, with the frequency-guided approach, Llama continues to outperform GPT-3.5 significantly, but both models’ performance metrics fall short of GPT-4o. This trend underscores the substantial performance gains achieved by combining the frequency-guided approach with GPT-4o, as discussed in Section 5.4.2.
7 Evaluation on Switchboard Dialog Act Corpus
To assess the generalizability of our approach, we evaluate two configurations on the SWBD-DAMSL annotation scheme (42 classes). The configurations are: (1) open-ended questions with a non-binary tree and (2) frequency-guided optimal split selection with backtracking. These configurations were selected as they represent the best non-frequency-guided and frequency-guided approaches. We use the resulting trees to annotate a randomly sampled subset of the Switchboard Dialogue Act Corpus Jurafsky et al. (1997b); Shriberg et al. (1998); Stolcke et al. (2000), [^10] GPL-2.0 license consisting of 260 utterances. [^11] We initially selected 300 utterances, but 40 were annotated as “+”, which, in this taxonomy, indicates that the utterance continues the label of the preceding one. Since these cases do not require actual annotation but rather a repetition of the previous label, we excluded them from the analysis.
Table 6 presents evaluation metrics using GPT-4o. The results confirm that the proposed framework is effective and generalizable across different taxonomies, including large-scale ones like SWBD-DAMSL, with the Frequency-Guided Optimal Split Selection achieving a Weighted F1 score of 0.61.
Approach $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
Non-binary 0.63 0.47 0.48 0.18
Freq.-guided split selection 0.65 0.63 0.61 0.23
Table 6: Evaluation of dialog act annotations from the SWBD-DAMSL annotation scheme, generated by GPT-4o, on a randomly selected set of 260 utterances from the Switchboard Dialogue Act Corpus using non-binary open-ended questions tree, and a frequency-guided optimal split selection tree.
8 Conclusions
We conducted experiments on generating tree schemes for discourse taxonomies using LLMs. This paper proposes a framework that supports the entire pipeline, from tree construction to dialog annotation. Our configuration with frequency-guided tree creation demonstrates that using LLMs for both tree scheme generation and annotation can yield results that surpass manual tree construction and crowdsourced annotation while significantly reducing the time required for the entire process.
Limitations
A key limitation of the proposed method is the restricted set of models that can be used. The tree creation process requires an advanced model, and the annotation step also benefits from using a more sophisticated model. Another limitation is that non-frequency-guided configurations still underperform compared to manually created trees. This highlights the importance of class frequency information in achieving optimal performance.
Potential directions for future research, motivated by the current limitations, include: (1) exploring larger open-source models for improved taxonomy generation and annotation; (2) conducting experiments on other domains, such as classroom discourse and task-oriented dialog systems; (3) incorporating human feedback to allow the model to self-correct and improve annotation accuracy; and (4) enabling self-refinement of the taxonomy by adapting to new, previously unseen dialog examples.
Ethical Considerations
Dialog data often contains personal or sensitive information, making it essential to anonymize and handle data securely when applying the proposed approach to individual datasets. This is crucial to safeguarding privacy rights. Beyond this consideration, we do not anticipate any significant risks associated with this work or the use of the proposed framework.
Appendix A Taxonomy of Speech Functions
Table 7: Taxonomy of speech functions (“-” indicates that these labels are counted together with Open.Command.)
Label Definition Example Frequency (%)
Open.Attend These are usually greetings. Hey, David! 1.6
Open.Demand.Fact Demanding factual information at the beginning of a conversation or when introducing a new topic. What’s Allenby doing these days? 2.7
Open.Demand.Opinion Demanding judgment or evaluative information from the interlocutor at the beginning of a conversation or when introducing a new topic. Do we need Allenby in this conversation? 1.1
Open.Give.Fact Providing factual information at the beginning of a conversation or when introducing a new topic. You met his sister. 1.8
Open.Give.Opinion Providing judgment or evaluative information at the beginning of a conversation or when introducing a new topic. This conversation needs Allenby. 0.9
Open.Command Making a request, an invitation or command to start a dialog or discussion of a new topic. Could you tell me about your wedding? 1.1
React.Rejoinder.Support.Track.Probe Requesting a confirmation of the information necessary to make clear the previous speaker’s statement. Because Roman lives in Denning Road also? 1.9
React.Rejoinder.Support.Track.Check Getting the previous speaker to repeat an element or the entire statement that the speaker has not heard or understood. Straight into the what? 0.9
React.Rejoinder.Support.Track.Clarify Asking a question to get additional information on the current topic of the conversation. Requesting to clarify the information already mentioned in the dialog. What, before bridge? 12.0
React.Rejoinder.Support.Track.Confirm Asking for a confirmation of the information received. [David: Well, he rang Roman, he rang Roman a week ago.]
Nick: Did he?
1.6
React.Rejoinder.Support.Response.Resolve The response provides the information requested in the question. [Fran: Oh what is it called?]
Brad: PhD in Science.
8.7
React.Rejoinder.Confront.Response.Re-challenge Offering an alternative position, often an interrogative sentence. [David: No, Messi is the best]
Nick: PAUSE
David: The best is Pele
0.2
React.Rejoinder.Confront.Challenge.Rebound Questioning the relevance or reliability of the previous statement, often an interrogative sentence. [David: This conversation needs Allenby.]
Fay: Oh he’s in London. So what can we do?
0.5
React.Rejoinder.Confront.Challenge.Detach Terminating the dialogue. So stick that! 0.5
React.Rejoinder.Confront.Challenge.Counter Dismissing the addressee’s right to his/her position. You don’t understand, Nick. 1.2
React.Rejoinder.Confront.Response.Refute Rejecting a transition to a new topic. [David: I’m out.]
Fay: You can’t do that, it’s my birthday.
0.1
React.Respond.Support.Register A manifestation of emotions or a display of attention to the interlocutor. Yeah. Right. Hmm… 6.0
React.Respond.Support.Engage Drawing attention or a response to a greeting. Hey! Hi-hi. 0.6
React.Respond.Support.Reply.Accept Expressing gratitude. Thank you! 1.2
React.Respond.Support.Reply.Affirm A positive answer to a question or confirmation of the information provided. Yes/its synonyms or affirmation. [Nick: He went to London.]
Fay: He did.
3.7
React.Respond.Support.Reply.Acknowledge Indicating knowledge or understanding of the information provided. I know. I see. Oh yea. 1.1
React.Respond.Support.Reply.Agree Agreement with the information provided. In most cases, the information that the speaker agrees with is new to him. Yes. Right. 3.8
React.Respond.Support.Develop.Extend Adding supplementary or contradictory information to the previous statement. David: [That’s what the cleaner—your cleaner lady cleaned my place though.]
Nick: She won’t come back to our place.
8.6
React.Respond.Support.Develop.Enhance Adding details to the previous statement, adding information about time, place, reason, etc. [Fay: He kept telling me I’ve got a big operation on with.]
Nick: The trouble with Roman though is that—you know he does still like cleaning up.
0.4
React.Respond.Support.Develop.Elaborate Clarifying/rephrasing the previous statement or giving examples to it. A declarative sentence or phrase (may include for example, I mean, like). [Nick: Cause all you’d get is him bloody raving on.]
Fay: He’s a bridge player, a naughty bridge player.
0.2
React.Respond.Confront.Reply.Disavow Denial of knowledge or understanding of information. I don’t know. No idea. 0.4
React.Respond.Confront.Reply.Disagree Negative answer to a question or denial of a statement. No, negative sentence. [David: Is he in London?]
Nick: No.
2.0
React.Respond.Confront.Reply.Contradict Refuting previous information. Sentence with opposite polarity. If the previous sentence is negative, then this sentence is positive, and vice versa. [Fay: Suppose he gives you a hard time, Nick?]
Nick: Oh I like David a lot.
0.4
React.Respond.Command Making a request, an invitation, or command in response to previous information. Could you tell me about your wedding? -
Sustain.Continue.Monitor Checking the involvement of the listener or trying to pass on the role of speaker to them. You know? Right? 0.2
Sustain.Continue.Command Making a request, an invitation, or command to continue the dialog or discussion without changing the speaker. Could you tell me about your wedding? -
Sustain.Continue.Prolong.Extend Adding supplementary or contradictory information to the previous statement. Used only when the speaker remains the same as in the previous utterance. Just making sure you don’t miss the boat. I put it out on Monday mornings. I hear them. I hate trucks. 21.8
Sustain.Continue.Prolong.Enhance Adding details to the previous statement, adding information about time, place, reason, etc. Used only when the speaker remains the same as in the previous utterance. Nor for much longer. We’re too messy for him. 5.1
Sustain.Continue.Prolong.Elaborate Clarifying/rephrasing the previous statement or giving examples to it. Used only when the speaker remains the same as in the previous utterance. Yeah but I don’t like people… um… I don’t want to be INVOLVED with people. 7.9
Appendix B Example of Tree Scheme
Figure 3: An example of a tree scheme generated by GPT-4 using our proposed approach, where nodes represent questions about an utterance, arrows indicate possible answer choices and purple leaves correspond to taxonomy labels.
Appendix C Example of dialog annotation with Speech Functions
Figure 4: Example of dialog annotation with speech functions Ostyakova et al. (2023).
Appendix D The Top Level Labels of SF taxonomy
Label Definition
Open Open utterances are statements or actions that initiate a conversation or introduce a new topic within an ongoing discussion. These may include greetings, questions, requests, invitations, or the sharing of information.
React React utterances are responses to the interlocutor’s statements. These may include answers to questions, follow-up questions, emotional reactions, sharing information, expressions of agreement or disagreement, and more.
Sustain Sustain utterances are those that extend the speaker’s own preceding statements by adding information, providing new details, or rephrasing. The “Sustain” label is applied only when the current and preceding utterances are made by the same speaker. These utterances cannot take the form of questions, except when the question serves to confirm that the listener is paying attention.
Table 8: Definitions of Open, React and Sustain labels written manually.
Appendix E The Top Two Level Classes of SF taxonomy
Label Definition
Open.Demand Questions at the beginning of a conversation or when introducing a new topic.
Open.Give Providing information or opinion at the beginning of a conversation or when introducing a new topic.
Open.Command Making a request, an invitation or command to start a dialog or discussion of a new topic.
Open.Attend These are usually greetings.
Sustain.Continue These are used only when there is no change in the speaker from the previous utterance, except for cases when the utterance is a reply to a greeting. The Sustain.Continue class involves adding additional information or details to the speaker’s previous statement. It can also include questions intended to check if the interlocutor is listening, as well as requests or invitations.
React.Rejoinder These include any type of question or detailed response to the interlocutor’s questions, expressions of emotion, and grounding utterances such as “hmm,” “aha,” and similar reactions.
React.Respond These include positive or negative responses to questions and expressions of understanding or misunderstanding. They also include the provision of new information or details, similar to Sustain.Continue. However, in this case, the speaker of the current utterance differs from the speaker of the previous one (this rule regarding speakers applies only to cases similar to Sustain.Continue).
Table 9: Definitions of the two highest-level labels.
Appendix F Example of the Model’s Reasoning Output
Figure 5: An example of question-answer pairs generated by GPT-4 to reason about identifying utterances of different classes.
Appendix G Prompt template for Tree Construction
Figure 6: The prompt template used for all experiments with tree construction.
Appendix H Prompt template for the annotation step
Figure 7: The prompt template used for the annotation step.
Appendix I Prompt for Constructing a Yes/No Questions Tree
Figure 8: The prompt used for constructing trees with yes/no questions.
Appendix J Prompt for Constructing an Open-Ended Questions Tree
Figure 9: The prompt used for constructing trees with open-ended questions.
Appendix K Prompt for Constructing a Non-Binary Open-Ended Questions Tree
Figure 10: The prompt used for constructing non-binary trees with open-ended questions.
Appendix L Prompt For Scoring Splits
Figure 11: The prompt used to score split-candidates.
Appendix M Prompt For Frequency-Guided Tree Creation
Figure 12: The prompt is designed to create a decision tree that enables splitting data into more than two groups and that specifically instructs the LLM to form one group that exclusively contains the single most frequent class